February field update by Kathy

Hello all,
Overview- The season is winding down.  The resupply ship is at the Pier and our Vessel samples have been staged to go north on the ship.  Vessel off-load is complete, but on-load is still underway.  The vessel is "early" again this year, so we had some beautiful days here early in the week, rather than the stormy weather that we normally associate with the arrival of the resupply vessel.  There is a strong southerly wind blowing today and sea-ice has been on the move all this week.  All of our team members have left the field and are based in McMurdo until we redeploy.  The remaining field groups are packing samples for COMAIR and finalizing close-out for the season.

C-504  Adam and Forrest have wrapped up their season and are back in McMurdo. 

C-505  Amy Chiuchiolo finished the remaining lab work for the limno team and left on Saturday. 

C-506 The stream team completed their stream monitoring this week.  They closed out the gage boxes for the winter and did end-of-season lake levels.  Aneliya is working on Chlorophyll analysis for Algae Ops.  The team is unpacking and sorting all of the retro.  They have 2 more day trips planned for this week. 

C-507 finished their work and went north earlier this week. 

C-509 spent a day sampling the Kaki Ponds, aka Marr Ponds, see photo attached.  The analytical team has been making good progress on lab work, but then had a series of issues with all of the analytical instruments.  As of 1/30/16 none of the instruments were working properly.  As of now, anion analyses are working, cations are still unknown and both Shimadzu instruments are having significant problems and we continue to work on troubleshooting.  We are working to resolve all of these issues and making contingency plans as we wrap up the season. 

I'll have two more updates before the end of the season. 


Article Section: